Evaluate the definite integral
∫728x−−√dx
1
Expert's answer
2020-12-07T16:08:31-0500

Evaluate the definite integral:

"\int^7_2\sqrt{8x} dx" = "2\sqrt{2}\int^7_2 \sqrt{x}dx=2\sqrt{2}(\frac{2\sqrt{x^3}}{3}\mid^7_2)=2\sqrt{2}(\frac{2\sqrt{7^3}}{3}-\frac{2\sqrt{2^3}}{3})"


"2\sqrt{2}(\frac{2\sqrt{7^3}}{3}-\frac{2\sqrt{2^3}}{3})=\frac{28\sqrt{14}-8}{3}"